OpenStudy (anonymous):

Compute the discriminant of the quadratic equation 10x^2+x-3=0

OpenStudy (aravindg):

\[\large \bf discriminant=b^2-4ac\]

OpenStudy (anonymous):

ok so \[1^{2}-4(10)(-3)\]

OpenStudy (aravindg):

where the quadratic equation is \[\large \bf ax^2+bx+c=0\]

OpenStudy (aravindg):

yep right!!

OpenStudy (anonymous):

1+120=121

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

thank you medal given